OpenStudy (samigupta8):

The pitch of a screw gauge is 0.5mm and there are 100 divisions on circular scale.The instrument reads +2 circular divisions when nothing is put in between its jaws.In measuring the diameter of a wire, there are 8 divisions on the main scale and 83rd circular scale division coincide with the reference line.Then the diameter of wire is

OpenStudy (samigupta8):

@parthkohli

OpenStudy (samigupta8):

I m getting a diff ans then theirs...

OpenStudy (samigupta8):

@priyar

OpenStudy (samigupta8):

My ans is 8.405 mm and they have given it as 4.405.. I don't know how they got that?

OpenStudy (samigupta8):

@imqwerty

imqwerty (imqwerty):

least count = pitch/(no. of divisions) total reading = main scale reading + (corrected circular scale reading x least count) \(8+\left[\left(83-2) \right) \times 0.005 \right]\) =8.405

OpenStudy (samigupta8):

That's my ans even...bt they have given something else..

imqwerty (imqwerty):

their answer is wrong :)

OpenStudy (priyar):

Yes your answer is correct! @samigupta8